A random sample of 150 recent donations at a certain blood bank reveals that 82 were type A blood. Does this suggest that the actual percentage of type A donations differs from 40%, the percentage of population having type A blood? Carry out a test of appropriate hypotheses using a significance level of 0.01. Would your conclusion have been different if a significance level of 0.05 has been used?
Answer #1
Concepts and reason

Sampling Distribution of Proportion:

The random sample of size n is taken from the population with sample proportion p^\hat p .

The sampling distribution of the sample proportion has mean pp and the standard deviation p(1p)n\sqrt {\frac{{p\left( {1 - p} \right)}}{n}} . Moreover, the sample proportion follows normal distribution for large sample size n.

The null and alternative hypothesis for proportion test is as shown below:

The null hypothesis is denoted as H0{H_0} and defined as,

H0:p=p0H0:pp0H0:pp0\begin{array}{l}\\{H_0}:p = {p_0}\\\\{H_0}:p \le {p_0}\\\\{H_0}:p \ge {p_0}\\\end{array}

Hence, pp be the population proportion and p0{p_0} is the hypothesized value.

The alternative hypothesis is denoted as H1{H_1} and defined as,

Ha:pp0(Twotailed)Ha:p<p0(Lefttailed)Ha:pp0(Righttailed)\begin{array}{l}\\{H_a}:p \ne {p_0}\left( {{\rm{Two tailed}}} \right)\\\\{H_a}:p < {p_0}\left( {{\rm{Left tailed}}} \right)\\\\{H_a}:p \ge {p_0}\left( {{\rm{Right tailed}}} \right)\\\end{array}

Hence, pp be the population proportion and p0{p_0} is the hypothesized value.

Fundamentals

The formula for sample proportion is,

p^=Xn\hat p\,\, = \,\,\frac{X}{n}

Formula for z-score of sample proportion is given below:

z=p^p0p0(1p0)nz = \frac{{\hat p - {p_0}}}{{\sqrt {\frac{{{p_0}\left( {1 - {p_0}} \right)}}{n}} }}

Where p is the population proportion and p^\hat p is the sample proportion.

Rejection region (two tailed):

Reject the null hypothesis if zzα/2orzzα/2z \ge {z_{\alpha /2}}{\rm{ or }}z \le - {z_{\alpha /2}} .

Rejection region (right tailed):

Reject the null hypothesis if zzα/2z \ge {z_{\alpha /2}} .

Rejection region (left tailed):

Reject the null hypothesis if zzα/2z \le - {z_{\alpha /2}} .

Let p be the population proportion of type A blood.

The null and alternative hypotheses are,

Null hypothesis: There is no difference in the actual percentage of the population having type A blood is 40 %.

H0:p=04{H_0}:p = \,\,0\cdot4

Alternative hypothesis: The actual percentage of the population having type A blood differs from 40 %.)

H:p #04

The specified level of significance is, α=0.01\alpha \, = \,0.01 .

From the standard normal distribution table, the critical values at 0.01 level of the significance for two tailed test is ±2.576 \pm 2.576

Rejection region: reject the null hypothesis if z2.576orz2.576z \ge 2.576{\rm{ or }}z \le - 2.576 .

Here, X is the number of observations, and n is the sample size.

Let the number of persons having a Type A blood be, X=82X = 82 .

Let the number of donations be, n=150n = 150 .

The sample proportion is,

p^=Xn=82150=0.5466\begin{array}{c}\\\hat p\,\, = \,\,\frac{X}{n}\\\\ = \frac{{82}}{{150}}\,\,\\\\ = 0.5466\\\end{array}

The formula for the test statistic is,

z=p^p0p0(1p0)nz = \frac{{\hat p - {p_0}}}{{\sqrt {\frac{{{p_0}\left( {1 - {p_0}} \right)}}{n}} }}

z=0.54660404×06150=3.66663.667\begin{array}{c}\\z = \frac{{0.5466 - 0\cdot4}}{{\sqrt {\frac{{0\cdot4 \times 0\cdot6}}{{150}}} }}\\\\ = 3.6666\\\\ \approx 3.667\\\end{array}

Since, the test statistic value of 3.667 exceeds the critical value of 2.576. So, reject the null hypothesis. It can be concluded that actual percentage of the population having type A blood differs significantly from 40%.

The specified level of significance is, α=0.05\alpha \, = 0.05 .

From the standard normal distribution table, the critical values at 0.05 level of the significance for two tailed test is ±1.96 \pm 1.96

Rejection region: reject the null hypothesis if z1.96orz1.96z \ge 1.96{\rm{ or }}z \le - 1.96 .

Since, the test statistic value of 3.667 exceeds the critical value of 1.96. So, reject the null hypothesis. It can be concluded that actual percentage of the population having type A blood differs significantly from 40%.

Ans:

The actual percentage of the population having type A blood differs significantly from 40%.

The actual percentage of the population having type A blood differs significantly from 40%.